# Just-in-time and distributed task representations in language models

### v1

@@ Abstract
Many of language models’ impressive capabilities originate from their in-context learning: based on instructions or examples, they can infer and perform new tasks without weight updates. In this work, we investigate when representations for new tasks are formed in language models, and how these representations change over the course of context. We focus on “transferrable” task representations—vector representations that can restore task context in another instance of the model, even without the full prompt. We show that these representations evolve in non-monotonic and sporadic ways, and are distinct from a more inert representation of high-level task categories that persists throughout the context. Specifically, models often condense multiple evidence into these transferrable task representations, which align well with the performance improvement based on more examples in the context. However, this accrual process exhibits strong locality along the sequence dimension, coming online only at certain tokens—despite task identity being reliably decodable throughout the context. Moreover, these local but transferrable task representations tend to capture minimal “task scopes”, such as a semantically-independent subtask, and models rely on more temporally-distributed representations to support longer and composite tasks. This two-fold locality (temporal and semantic) underscores a kind of just-in-time computational process underlying language models’ ability to adapt to new evidence and learn new tasks on the fly.
@@ 1 Introduction
Much of the excitement about large language models began with the discovery that they exhibit In-Context Learning (ICL; Brown et al.,, 2020): the emergent ability to learn tasks from few-shot examples in context. This discovery has led to a variety of works exploring the behavioral features of ICL (e.g. Sclar et al.,, 2024; Min et al.,, 2022). Other works have studied the dynamics of ICL, and how performance improves with increasing numbers of few-shot examples (Agarwal et al.,, 2024; Anil et al.,, 2024). The strong behavioral success of ICL led to substantial interest in understanding the mechanistic basis of these capabilities. Work on interpreting the mechanisms responsible for ICL has led to discoveries such as induction heads (e.g. Olsson et al.,, 2022) and how ICL implicitly refines a model of in-context evidence (e.g. Akyürek et al.,, 2022; Von Oswald et al.,, 2023).
@@
Recently, several works have identified internal, vector-form task representations that can be extracted from a model’s forward pass on a few-shot prompt (Todd et al.,, 2024; Hendel et al.,, 2023). Importantly, these task representations not only capture general task information, but can be used to restore the appropriate task context during the model’s forward pass on a zero-shot prompt. This transfer effect is observed by intervening with that representation at the appropriate place in the model’s residual stream—such that task context is reinstantiated and the model can perform the task without any explicit instructions or demonstrations. These “transferrable” task representations have been shown to exist across a variety of tasks and presentation formats, and even capture transferrable task knowledge across modalities (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024).
@@
The discovery of transferrable task representations raises several intriguing questions about their dynamics and generality. How and when are these internal task representations formed throughout the context? How do these dynamics depend upon task complexity? A simple, intuitive hypothesis is that these representations develop gradually, e.g. at a rate that depends on task complexity. The representations might accrue evidence from each example and refine monotonically into a more stable, robust task representation. This view aligns with the behavioral findings that models perform better with more examples in-context (Agarwal et al.,, 2024; Anil et al.,, 2024).
@@
We set out to understand how the dynamics of ICL (within and across examples) are exhibited representationally within the model, by investigating task representations. Our findings suggest a more complicated picture of the computational process that support language models’ adaptation to a new task:
@@
  • For simple tasks, we find two types of task representations in language models: an inert representation of task identity is more continuously present throughout the context, but transferrable task representations only activate sporadically at key tokens.
@@
  • These fleeting but transferrable task representations often condense evidence from multiple examples but tend to capture minimal task scopes, as their ability to guide model behavior decays over longer generation and across independent subtask contexts.
@@
Figure 1: Understanding how task representations develop over context. A. A schematic of extracting transferrable task representations and restoring task contexts in zero-shot settings. The highlighted tokens indicate the source and target for extracting and injecting task representations. B. Transferrable task representations restore task accuracy on zero-shot prompts. Results are aggregated over all models for simple tasks (see Appendix A). Error bars indicate the 95% CI over tasks. C. An overview of the development of transferrable task representations over context, i.e. the recontextualized zero-shot accuracy for task representations extracted from different tokens throughout the context.
@@ 2 Methods
For our analyses, we built upon tasks from prior work on transferrable task representations (Hendel et al.,, 2023; Todd et al.,, 2024). The tasks we examine include a diverse set of natural language tasks (e.g., finding the antonym of a query word, translating an English word to French) and algorithmic tasks (e.g., counting or extracting a target word from a list of input words). In addition to these simple, single-token generation tasks from the previous literature, we also test a range of new tasks to explore model behavior in longer generation settings. These include: repeating a simple task three times (e.g. antonym x 3 requires finding the antonyms of three input words), extracting multiple words from a query word list (e.g., choose the first and last words from a word list), and reversing or shifting an entire word list. Finally, we also explore a set of “mixed-generation” tasks, where the model needs to infer and perform different tasks on each word in a word list. (See Appendix A for the full set of tasks.)
There are 512 query-answer pairs for each task (except for two smaller datasets: country-capital contains 197 samples, and product-company contains 494 samples). These query-answer pairs are formatted into few-shot prompts with alternating "Q:" and "A:" turns, as shown in Figure 1A.
@@
We focus our analyses on the open-weight Gemma V3 pre-trained models, including the 4B, 12B, and 27B-sized models (Team et al.,, 2025).
Extracting in-context task representations
@@
We primarily investigate task vectors discovered in (Hendel et al.,, 2023) as a window to study language models’ in-context task representations. We extract task vectors from few-shot prompts consisting of query-answer pairs and a test query as shown in Figure 1A. Task vectors are the layer residual activations extracted from the last token before answer generation in the few-shot prompts (in the example in Figure 1A, this corresponds to the highlighted colon). Hendel et al., (2023) showed that task vectors can reinstate task performance even without any additional context and with a different query—when task vectors are patched onto (i.e., overwrite) the layer residual activations of the last token, they can recontextualize the model with the appropriate task context and enable the model to generate the task output without any few-shot examples in-context.
@@
We replicate and extend the procedure outlined in Hendel et al., (2023). For each model and task, we first search for the layer that best captures the task representation, using 50 queries from the dataset as the development set and in an 8-shot setting. As in prior work, we replace the real test queries with dummy queries sampled from the dataset to extract query-agnostic, general task representations. We searched among every 3 layers starting at layer 2 (0-indexed) for the 4B and 12B models (covering both the local-attention layers and global attention layers in Gemma V3 models; Team et al.,, 2025), and every 6 layers starting from layer 5 in the 27B model. The layer that restores the highest task accuracy on zero-shot prompts in the development set is designated as the layer that best captures the representation for a given task, and subsequently used to extract task vectors and restore task contexts for the remainder, held-out queries in the dataset. Consistent with prior results, we generally find that task vectors extracted and injected at middle layers restore the highest task accuracy on zero-shot prompts, for all model sizes.
Evaluating task accuracy
@@
We compare the average accuracy for the sampled responses across a few settings: standard zero-shot, recontextualized zero-shot setting (with task vector intervention), and few-shot. For simplicity, responses for all tasks are graded by exact string matches against the ground-truth answer. This underestimates the model performance in some tasks (e.g. for antonym and translation tasks), but we use the same grading scheme across all settings and compare relative performances. For longer-generation and mixed-generation tasks, we evaluate each of the multiple outputs separately by exact match (e.g., in antonym x 3, we compare each of the three output words with the correct answer), and report the mean accuracy across all outputs in the longer response.
Examining the dynamics of task representations
@@
Once we determine the best layer for each task using the last colon token, we evaluate how well the colon token representations condense the k-shot information in a prompt. We do this by extracting task vectors at the colon token for a k-shot prompt, and then evaluating the recontextualized zero-shot accuracy when the task vectors are patched into a zero-shot prompt. We repeat this analysis for k in 1, 2, 4, 8, 16, 32. We also experimented with allowing the best layer to vary depending on k, but found very similar results.
@@
We also studied whether transferrable task representations form in tokens other than the final colon token. In particular, we extracted layer residual token activations for other tokens in the context, including the "Q", the ":" following "Q", the "A", and the new-line token before "A". We patched these token activations onto the corresponding token in the zero-shot prompt at the same layer. For each of the non-final tokens, we repeated the search for the layer that best captures task representations. All token representations were evaluated on the extent to which they restored task accuracy on zero-shot prompts.
@@ 3 Results
How do language models infer new tasks in-context? We leverage the ability for in-context task representations to restore task contexts to understand how models accrue evidence and refine task representations. We find that language models indeed form stronger task representations that aggregates in-context evidence. However, this evidence accrual process is surprisingly non-gradual and happens in a sporadic way for most tasks (Figure 1C). In particular, effective and transferrable task representations form only at certain tokens in the context, and tend to capture the minimal “task scope.” For some tasks, models also do not appear to condense task information into local task representations. Below, we discuss these findings in more detail.
Figure 2: Sporadic & inconsistent evidence accrual in language models. A. Task vectors extracted from the last colon token in each example capture evidence accrual on most tasks (11 out of 14). However, on three “hard-to-transfer” tasks, task vectors do not capture this evidence accumulation, even though the models (behaviorally) do learn from more examples. The solid bars indicate recontextualized zero-shot accuracy (via task vectors), and light bars in the background indicate 8-shot accuracy (without task vectors). The dotted lines indicate the ratio of the recontextualized zero-shot accuracy against 8-shot accuracy. B. Most other format tokens in the context do not robustly form transferrable task representations that support recontextualization on zero-shot, but task identity is reliably decodable in their residual activations. Here, we report the task identity decoding accuracy at the mode best layer at which transferrable task representations form in the second ":" token. See the main text and Appendix C for more details.
@@ 3.1 Local task representations can accrue evidence
Task representations can reflect evidence accrual.
Consistent with the behavioral performance gain from including more examples in-context (e.g., Anil et al.,, 2024), we find that transferrable task representations in models also reflect increased task certainty with increased context. As shown in Figure 2A, task vectors extracted from more examples are better at restoring task performance in zero-shot settings, such that the ratio between the recontextualized zero-shot accuracy and few-shot accuracy stays relatively stable across the number of examples. This suggests that language models condense information from multiple examples and form better task representations, even when the task representations extracted are fairly local (i.e., at the last token in the prompt).
@@
However, we did not observe evidence accrual in 3 out of 14 of the simple tasks (Figure 2A, hard-to-transfer tasks). Specifically, for choose_middle_of_5, count_color_in_3, and count_fruit_in_3 tasks, we found that local task representations (extracted from the last token) were not able to take advantage of more examples for task transfer to zero-shot settings, even though models improved substantially at solving the task when given more examples in the prompt. One interpretation is that some tasks require more state-tracking (such as counting), and that these may necessitate additional inference processes that the models do not condense into local task representations. Alternatively, these inference processes cannot be effectively re-activated by the injection of the extracted task representations.
How evidence accrual leads task representations to converge.
@@
For the eleven other simple tasks where we successfully observed evidence accrual, we sought to understand how the task representations themselves changed over more examples (Figure 3A). As we increased the number of examples in context, we generally found reduced variance among task vectors extracted from different k-shot samples. This suggests that, as models gain evidence, in-context task representations tend to denoise or converge to more stable representations. The magnitude (L2-norm) of the task vectors also tends to decrease over time. However, there were noticeable differences between tasks and models. Some tasks seem to converge to stable representations faster (i.e., with fewer examples; see also a visualization of the representational trajectories in Figure S2). However, for certain tasks, the magnitude of the task representations first increases then decreases given more examples. We note that for this analysis, we look at the task representations extracted from the mode best layer across different tasks. This is to control for magnitude difference of the residual activations across layers and make a fair comparison. Although the best layer for transferrable task representations does sometimes differ across tasks, the best layer across tasks tend to reside in the middle layer range across all model sizes, consistent with prior findings (Hendel et al.,, 2023).
@@
Figure 3: Analyses of extracted task representations. A. The extracted task vectors (at the last colon token) decrease in both variance and magnitude with more examples, exhibiting a general tendency to condense evidence and converge onto stable task representations. The solid line shows the average across tasks, the transparent lines show the individual tasks. B. Extracted task vectors from the 27B model form distinct clusters. The numbers label the centroid for each task (see legend and results for other models in Figure S4). Task vectors are similar but distinguishable when a task was evaluated independently vs. embedded within a larger task structure, e.g., representations for antonym (0), antonym x 3 (14), and where antonym appears as a first task in a mixed-generation task chain (24&27) are close but distinct.

### v2

@@ Abstract
Many of language models’ impressive capabilities originate from their in-context learning: based on instructions or examples, they can infer and perform new tasks without weight updates. In this work, we investigate when representations for new tasks are formed in language models, and how these representations change over the course of context. We focus on “transferrable” task representations—vector representations that can restore task contexts in another instance of the model, even without the full prompt. We show that these representations evolve in non-monotonic and sporadic ways, and are distinct from a more inert representation of high-level task categories that persists throughout the context. Specifically, when more examples are provided in the context, transferrable task representations successfully condense evidence. This allows better transfer of task contexts and aligns well with the performance improvement. However, this evidence accrual process exhibits strong locality along the sequence dimension, coming online only at certain tokens—despite task identity being reliably decodable throughout the context. Moreover, these local but transferrable task representations tend to capture minimal “task scopes”, such as a semantically-independent subtask. For longer and composite tasks, models rely on more temporally-distributed representations. This two-fold locality (temporal and semantic) underscores a kind of just-in-time computational process that language models use to perform new tasks on the fly.
@@ 1 Introduction
Much of the excitement about large language models began with the discovery that they exhibit In-Context Learning (ICL; Brown et al.,, 2020): the emergent ability to learn tasks from few-shot examples in context. This discovery has led to a variety of works exploring the behavioral features of ICL (e.g. Sclar et al.,, 2024; Min et al.,, 2022). Other works have studied the dynamics of ICL, and how performance improves with increasing numbers of few-shot examples (Agarwal et al.,, 2024; Anil et al.,, 2024). The strong behavioral success of ICL led to substantial interest in understanding the mechanistic basis of these capabilities, leading to discoveries such as induction heads (e.g. Olsson et al.,, 2022) and how ICL implicitly refines a model of in-context evidence (e.g. Akyürek et al.,, 2022; Von Oswald et al.,, 2023).
@@
Recently, several works have identified internal, vector-form task representations that can be extracted from a model’s forward pass on a few-shot prompt (Todd et al.,, 2024; Hendel et al.,, 2023). These task representations not only capture general task information, but can be used to restore the appropriate task context during the model’s forward pass on a zero-shot prompt. This transfer effect is observed by intervening with that representation at the appropriate place in the model’s residual stream. Such intervention reinstates the task context and allows the model to perform the task without any explicit instructions or demonstrations. These “transferrable” task representations have been shown to exist across a variety of tasks and presentation formats, and even capture transferrable task knowledge across modalities (Davidson et al.,, 2025; Huang et al.,, 2024; Luo et al.,, 2024).
@@
The discovery of transferrable task representations raises several intriguing questions about their dynamics and generality. How and when are these internal task representations formed throughout the context? How do these dynamics depend upon task complexity? A simple, intuitive hypothesis is that these representations develop gradually, e.g. at a rate that depends on task complexity. The representations might accrue evidence across tokens and refine monotonically into a more stable and robust task representation. This view aligns with the behavioral findings that models perform better with more examples in-context (Agarwal et al.,, 2024; Anil et al.,, 2024).
@@
We set out to investigate these transferrable task representations to understand how the dynamics of ICL are exhibited within them. Our findings suggest a more complicated picture of the computational process that support language models’ adaptation to a new task:
@@
  • We find two types of task representations in language models: an inert representation of task identity is more continuously present throughout the context, but transferrable task representations only activate sporadically at key tokens.
@@
  • The fleeting but transferrable task representations condense evidence from multiple examples. However, they also tend to capture minimal task scopes, as their ability to guide model behavior decays over longer generation and across independent subtask contexts.
@@
Figure 1: Understanding how task representations develop over context. A. A schematic of extracting transferrable task representations and restoring task contexts in zero-shot settings. The highlighted tokens indicate the source and target for extracting and injecting task representations. B. Transferrable task representations restore task accuracy on zero-shot prompts. Results are aggregated over all models for simple tasks (see Appendix A). Error bars indicate the 95% CI over tasks. C. An overview of the development of different task representations over context. Solid bars: recontextualized zero-shot accuracy for task vectors extracted from different tokens. Transparent bars: task identity decoding accuracy from different token representations.
@@ 2 Methods
For our analyses, we built upon tasks from prior work on transferrable task representations (Hendel et al.,, 2023; Todd et al.,, 2024). The tasks we examine include a diverse set of natural language tasks (e.g., finding the antonym of a query word or translating an English word to French) and algorithmic tasks (e.g., counting or extracting a target word from a list of input words). In addition to these simple, single-token generation tasks from the previous literature, we also test a range of new tasks to explore model behavior in longer generation settings. These include: repeating a simple task three times (e.g. antonym x 3 requires finding the antonyms of three input words), extracting multiple words from a query word list (e.g., choose both the first and the last word in the list), and reversing or shifting an entire word list. Finally, we also explore a set of “mixed-generation” tasks, where the model needs to infer and perform different tasks on each input item. See Appendix A for the full set of tasks.
There are 512 query-answer pairs for each task (except for two smaller datasets: country-capital contains 197 samples, and product-company contains 494 samples). These query-answer pairs are formatted into few-shot prompts with alternating “Q:” and “A:” turns, as shown in Figure 1A.
@@
In the main paper, we present results on the open-weight pre-trained 4B, 12B, and 27B Gemma V3 models (Team et al.,, 2025). In Appendix C, we also show that the main findings replicate on the 4B, 8B, and 14B Qwen3 models (Yang et al., 2025a, ).
Extracting transferrable task representations
@@
We primarily investigate task vectors discovered in Hendel et al., (2023) as a window to study language models’ transferrable task representations. We extract task vectors from few-shot prompts consisting of query-answer pairs and a test query, as shown in Figure 1A. Task vectors are the layer residual activations extracted from the last token before answer generation (in the example in Figure 1A, this corresponds to the highlighted colon token). Hendel et al., (2023) showed that task vectors can reinstate task performance on a different query even without any prior context. Specifically, when task vectors are patched onto (i.e., overwrite) the layer residual activations of the last token, they can recontextualize the model with the appropriate task context and enable the model to generate the task output without any prior few-shot examples.
@@
We replicate and extend the procedure outlined in Hendel et al., (2023). For each model and task, we first search for the layer that best captures the task representation, using 50 queries from the dataset as the development set and in an 8-shot setting. As in prior work, we replace the real test queries with dummy queries sampled from the dataset to extract query-agnostic, general task representations. We searched among every 3 layers starting at layer 2 (0-indexed) for the 4B and 12B models (covering both the local-attention layers and global attention layers in Gemma V3 models; Team et al.,, 2025), and every 6 layers starting from layer 5 in the 27B model (covering the global attention layers). The layer that restores the highest task accuracy on zero-shot prompts in the development set is designated as the layer that best captures the representation for a given task. This best layer is subsequently used to extract task vectors and restore task contexts for the remainder, held-out queries in the dataset. Consistent with prior results, we generally find that task vectors extracted and injected at middle layers restore the highest task accuracy on zero-shot prompts, for all model sizes.
Evaluating task transfer
@@
We compare the average accuracy of the sampled responses across three settings: standard zero-shot, recontextualized zero-shot (with task vector intervention), and few-shot (with examples in context). For simplicity, responses for all tasks are graded by exact string matches against the ground-truth answer. This underestimates the model performance in some tasks (e.g. for antonym and translation tasks), but we use the same grading scheme across all settings and compare relative performances. For longer-generation and mixed-generation tasks, we evaluate each of the multiple outputs separately by exact match (e.g., in antonym x 3, we compare each of the three output words with the correct answer), and report the average accuracy across all output units.
Examining the dynamics of transferrable task representations
@@
Once we determine the best layer for each task using the last colon token, we evaluate how well the colon token representations condense information from multiple examples in a prompt. We do this by extracting task vectors at the colon token in different k-shot prompts and patching onto zero-shot prompts, then comparing the recontextualized zero-shot accuracy. We repeat this analysis for k in 0, 1, 2, 4, 8, 16, 32. In an earlier experiment on a subset of the tasks, we also experimented with allowing the best layer to vary depending on k, but found very similar results overall. We show these layer search results across different k’s in Figure 2A, but otherwise focus on results from reusing the best layer from the layer search on 8-shot prompts for other k’s.
@@
We then studied whether transferrable task representations form in tokens other than the colon token. We extracted layer residual token activations for other format tokens in the context, including the “Q”, the “:” following “Q”, the “A”, and the new-line token before “A”. We patched these token activations onto the corresponding token in the zero-shot prompt at the same layer. For each of the non-colon tokens, we repeat the search for the layer that best captures task representations. All token representations are evaluated on the extent to which they restore task accuracy on zero-shot prompts.
Decoding task identity
We also studied whether token representations throughout the context contain robust task identity information. We trained simple linear decoders to predict the task category from the layer residual activations of different tokens. At each layer and token combination, we used 100 token instances for each task to train and test a task identity decoder. All decoders were trained for 20 epochs, with a batch size of 256 and a learning rate of 0.01. We report the decoding accuracy across the 25% held-out test representations.
@@ 3 Results
How do language models infer new tasks in-context? We leverage the transferability of task contexts to understand how models accrue evidence and refine task representations. We find that language models indeed form stronger task representations that aggregate in-context evidence. However, this evidence accrual process is non-gradual and happens in a sporadic way. In particular, models only form effective task representations that transfer task contexts at certain tokens (Figure 1C, also see Figure 2A). This is in strong contrast to a persistent sensitivity of task identity across tokens in the context (Figure 2B). We also find that the surprisingly local transferrable task representations tend to only capture a minimal “task scope.” Below, we discuss these findings in more detail.
Figure 2: Transferrable task representations activate sporadically at key tokens, but task identity representations persist throughout the context. A. Recontextualization accuracy when each token representation is used to restore task contexts in zero-shot settings. B. Task identity decoding accuracy (among 14 tasks) for token representations at different layers and positions. This figure plots aligned sequences across different samples and tasks; since exact positions differ depending on the sample, the indices shown in the labels are approximate. See results for other models in Figure S4.
Figure 3: Sporadic & inconsistent evidence accrual in language models. A. Task vectors extracted from the last colon token in each example capture evidence accrual on most tasks (12 out of 14). However, on two “hard-to-transfer” tasks, task vectors do not capture this evidence accumulation, even though the models (behaviorally) do learn from more examples. The solid bars indicate recontextualized zero-shot accuracy (via task vectors), and light bars in the background indicate few-shot accuracy (without task vectors). The dotted lines indicate the ratio of the recontextualized zero-shot accuracy against few-shot accuracy. B. Most other format tokens in the context do not robustly form transferrable task representations that support recontextualization on zero-shot, but task identity is reliably decodable in their residual activations. Here, we report the task identity decoding accuracy at the mode best layer at which transferrable task representations form in the second “:” token. See the main text for more details.
@@ 3.1 Local task representations can accrue evidence
Transferrable task representations reflect evidence accrual.
Consistent with the behavioral gain from including more examples in-context (e.g., Anil et al.,, 2024), we find that transferrable task representations also reflect increased task certainty with increased context (Figure 3A). Task vectors extracted following more examples are better at restoring task performance in zero-shot settings, such that the ratio between the recontextualized zero-shot accuracy and few-shot accuracy stays relatively stable across the number of examples. This suggests that language models condense information from multiple examples and form better task representations, even though the task representations extracted are fairly local (i.e., from a single token in the few-shot prompts).
@@
However, we did not observe strong evidence accrual in 2 out of the 14 simple tasks (Figure 3A, hard-to-transfer tasks). These two tasks are count_color_in_3 and count_fruit_in_3. For all three model sizes, task vectors extracted from 32-shot prompts recovered below 50% of the few-shot accuracy with 32 examples in the context. In other words, local task representations for these tasks were not able to take advantage of more examples for task transfer to zero-shot settings, even though models improved substantially at solving the task when given more examples in the prompt. Interestingly, the same counting tasks were hard-to-transfer for Qwen3 models as well (see Figure S7A). One possibility is that these tasks require more state-tracking, which may necessitate additional inference processes that the models do not condense into local task representations. Alternatively, these inference processes cannot be effectively re-activated by the injection of the extracted task representations.
How evidence accrual leads transferrable task representations to converge.
@@
As models appear to successfully accrue evidence in these highly local representations, we sought to understand how the task representations themselves changed over more examples (Figure 4A). For this analysis, we look at the task representations extracted from the mode best layer across different tasks. This is to control for magnitude differences of the residual activations across layers and make a fair comparison. Although the best layer for transferrable task representations sometimes differ across tasks, the best layers tend to reside in the middle layer range across all model sizes, consistent with prior findings (Hendel et al.,, 2023).
In Gemma V3 models, as we increase the number of examples in context, we generally found reduced variance among task vectors extracted from different k-shot prompts. This can signal that in-context task representations tend to denoise or converge to more stable representations as models gain evidence. The magnitude (L2-norm) of the task vectors also tends to decrease over time. However, for both the variance and magnitude, there were considerable differences between tasks and model sizes. Some tasks seem to converge to stable representations faster (i.e., with fewer examples; see also a visualization of the representational trajectories in Figure S2). For certain tasks, the magnitude of the task representations first increases then decreases given more examples. We also note that both the variance and magnitude of task vectors tend to increase per more examples in some Qwen3 models (Figure S8), suggesting that different models may develop different strategies for refining task representations.
@@
Figure 4: Analyses of extracted task representations in Gemma V3 models. A. The extracted task vectors (at the last colon token) tend to decrease in both variance and magnitude with more examples, exhibiting a general tendency to condense evidence and converge onto stable task representations. The solid line shows the average across tasks. The transparent lines show the individual tasks. B. Extracted task vectors from the 27B model form distinct clusters. The numbers label the centroid for each task (see legend and results for other models in Figure S5). Task vectors are similar but distinguishable when a task is evaluated independently vs. embedded within a larger task structure. For example, representations for antonym (0), antonym x 3 (14), and where antonym appears as a first task in a mixed-generation task chain (24&27) are close but distinct.